OpenStudy (firejay5):

Solve each system of equation by elimination: Show work and steps 21. 3x - 5y = 6 2x - 4y = 4

OpenStudy (anonymous):

2(3x-5y)= 2x6 6x-10y=12 -3(2x-4y)=-3x4 -6x-12y=-12 solve the above 2 equations and you get y=0 , substitute the value of y in the first equation and you will get x=2

OpenStudy (firejay5):

thanks! @vm12 you're the best! :D

OpenStudy (anonymous):



OpenStudy (firejay5):

@RaphaelFilgueiras I already had help thank you anyway

OpenStudy (anonymous):

ok yw